# Uncertainty in Graph Neural Networks: A Survey

### v1

@@ 3.3 Other Methods
Ensemble Model. Different from the Bayesian Gnns that perturb parameters around a single optimal model, ensemble models derive the final prediction based on the predictions from a set of independent models with different parameters or even with different model structures. Ensemble Gnns have been shown to provide more accurate prediction and reliable estimation than the single models (bazhenov2022towards, ; wang2023trafficuncertainty, ; gawlikowski2023survey, ). Meanwhile, ensemble Gnns offer an intuitive means of expressing predictive uncertainty by assessing the diversity among the predictions of the ensemble members. AU and EU could also be differentiated similarly as in Bayesian Gnn methods. Despite their wide applications, most Gnn ensemble models follow the deep ensemble framework (lakshminarayanan2017simple, ) without specific modifications for both graphs and uncertainty quantification purposes. Compared with the two families above, ensemble methods are less used in Gnns mainly due to their high computational cost.
Test-time data augmentation, another type of uncertainty quantification method, has been widely used in other fields, e.g., computer vision (ayhan2022test, ). It involves augmenting multiple samples for each test sample and applying the model to all samples to compute the distribution of predictions. Then, AU is estimated with the entropy of the model prediction distribution for a test sample and its augmentations, and EU can also be estimated in the same way as Bayesian Gnn methods. However, in Gnns, test-time augmentation has only been utilized for other purposes. jin2022empowering first propose to modify the node feature and the graph topology at the test time for better OOD generalization. Similarly, ju2024graphpatcher use test time augmentation to mitigate the bias of Gnn towards achieving better performance on high-degree nodes. However, test-time augmentation has not yet been explicitly applied for graph uncertainty quantification and thus, is not included in our primary framework. Nevertheless, its potential adaptation for graph uncertainty quantification warrants exploration. Given its post hoc nature, it requires no additional data and shows better computational efficiency during training compared to ensemble and Bayesian Gnns.
@@
3.4 Evaluating Uncertainty Quantification Models
@@
We discuss the evaluation measures to assess the quality of Gnn uncertainty quantification models. As in Figure 2, evaluations are not limited to a specific type of quantification model but to each targeted uncertainty source. This is because a model can quantify multiple sources of uncertainty, and models can only be evaluated regarding a specific uncertainty source. Given the challenge of acquiring ground truth uncertainty, various metrics have been proposed, and there is no consensus on a unified metric (lakshminarayanan2017simple, ; gawlikowski2023survey, ).
Predictive (Total) uncertainty. As discussed above (liu2022confidence, ; zhao2020uncertainty, ), maximum softmax probability (referred to as confidence in this paper) and entropy are widely applied to quantify the predictive uncertainty in the classification problems. They are also utilized as uncertainty measures of predictive uncertainty smith2018understanding . Notably, as they are affected by both aleatoric and epistemic uncertainty, we do not classify them into data uncertainty evaluation metrics as in gawlikowski2023survey . In regression problems, the mean width of prediction intervals (MWPI) (kang2022jurygcn, ; huang2023uncertainty, ) can be used for evaluation. The coverage rate—the portion of ground-truth labels falling into these intervals—is verified to ensure the validity of this metric. MWPI usually estimates the predictive uncertainty in all data instead of individual samples, and can only provide prediction intervals with uniform probability mass. These metrics are vulnerable to distribution shifts and need calibration.
@@
Model & Data uncertainty. The model uncertainty is commonly measured through mutual information (MI) to evaluate quantification models. MI is minimal when information in prediction does not increase with additional model knowledge. It is calculated as the difference between the entropy of the expected distribution (predictive uncertainty) and the expected entropy (data uncertainty) (malinin2018predictive, ; zhao2020uncertainty, ). Notably, MI is also referred to as information gain in some papers (liu2022confidence, ). For Bayesian and ensemble Gnns, the predictive variance of sampled outputs is another option to interpret model uncertainty (zhang2019bayesianchemical, ; gal2016dropout, ). Furthermore, employed as one of the optimization objectives in some quantification methods (liu2020uncertainty, ; zhao2020uncertainty, ), the expected Kullback-Leibler (KL) divergence between approximated distribution and the posterior distribution can also evaluate model uncertainty. Given no distributional uncertainty, model uncertainty equals epistemic uncertainty. In this case, data uncertainty, namely aleatoric uncertainty, can be measured as the difference between the total uncertainty and model uncertainty; otherwise, the metric will be no longer meaningful.
@@
Distributional Uncertainty. It is believed that quantification of distributional uncertainty will be able to separate in-distribution (ID) and out-of-distribution (OOD) samples malinin2018predictive . A popular approach to measure total distributional uncertainty is through classifying test samples. Samples with distributional uncertainty surpassing a given threshold are categorized as OOD samples, while those below the threshold are considered ID samples. In this, conventional classification metrics, e.g., Receiver Operating Characteristic (ROC) curve and Area Under Receiver Operating Curve (AUROC) (zhao2020uncertainty, ; song2022learning, ; gui2022good, ), can be applied to measure the distributional uncertainty. However, this evaluation depends on the selection of test samples and threshold.
In conclusion, the effectiveness of existing metrics remains unclear. Therefore, to derive reliable and straightforward measures, a wide range of papers directly apply specific downstream tasks for evaluation.
@@ 4.1 Uncertainty-based Node Selection
Active Learning. In graph active learning tasks, uncertainty and representativeness (diversity) are prevalent criteria for node selection  (ren2021survey, ). Uncertainty is often measured by the entropy of the prediction probabilities on the nodes. cai2017active first propose a graph active learning framework, which uses a combination of entropy-measured uncertainty, information density, and node centrality to select nodes. zhang2021alg estimate an ensembled uncertainty using the variation of prediction generated by committee models, which is more robust and reliable than entropy as a single deterministic model. The method selects top $k$ nodes to maximize the effective reception field using the topological information of the graph, excluding nodes that are less uncertain and less informative. Unlike the previous two methods, which select the most uncertain nodes independently, zhang2021information consider label propagation and select a subset of nodes with maximized uncertainty (entropy) reduction to include diversity. Labeled nodes can propagate their label information based on the graph distance and the node feature similarity, and then reduce the uncertainty of adjacent unlabeled nodes. Additionally, in a multi-task manner, chang2024multitask estimate the confidence difference between the node classification task and anomaly detection task for node selection in graph active learning. Notably, in most cases, graph active learning methods that select nodes with only uncertainty can hardly outperform the ones with a combination of uncertainty and other measures.
@@
To better label the nodes with dynamic and adaptive criteria, reinforcement learning (RL) has been applied to graph active learning. gao2018active use the combination of the measures in cai2017active to select nodes, but dynamically adjusts the combination weights based on a multi-armed bandit framework. Instead of measuring nodes independently, hu2020graph formulate the active learning problem as a sequential decision process on the graph considering the interconnections between nodes. Uncertainty is still measured with entropy but concatenated with other features, e.g., the divergence between a node’s predicted label distribution and its neighbor’s, to form the state representation of each node. The reward is the model performance on hold-out data updated with a trajectory of selected nodes. However, the RL-based active learning methods are significantly time-consuming to train.
@@
Self-training. In graph self-training, i.e., pseudo-labeling, li2018deeper propose one of the initial works that expand the labeled set with top $K$ confident nodes, quantified through maximal softmax probabilities. Considering that Gnn is ineffective in propagating label information under the few-label setting, sun2020multi iteratively select and label the top $K$ confident nodes. Drawing upon the concept of ensembling to robustly select and label nodes, yang2021self employ multiple diverse Gnns. Nodes are selected with confidence surpassing a pre-determined threshold and ensuring consistent prediction for labels across all models. wang2021confident is the first work to point out that the under-confidence of Gnns leads to biased uncertainty estimation, deteriorating the performance of uncertainty-based self-training node selection. With the calibrated confidence, they improve the performance of the self-training method introduced by li2018deeper . In consideration of graph structure information, zhao2021entropy propose to compute entropy for each node, and add an entropy-aggregation layer to Gnn for uncertainty estimation. Based on an expansion assumption (wei2021theoretical, ), wang2023deep prove that graph self-training error is bounded by the confidence threshold of the top $K$ confident nodes and the consistency of multi-view predictions. Therefore, they select nodes with the highest average confidence from multiple diversely augmented predictors.
@@
Similar to graph active learning, recent self-training methods (liu2022confidence, ; li2023informative, ) consider combining uncertainty with representativeness, but from different perspectives. liu2022confidence believe the common self-training practice that only selects high-confidence nodes leads to the distribution shift from the original labeled node distribution to the expanded node distribution. Therefore, they follow the common practice that selects high-confidence nodes as pseudo-labels but decreases the weight of the nodes with low information gain in the training loss. Interestingly, the dropout technique employed to compute information gain is often interpreted as the model uncertainty, as introduced in Section 3.4. Similarly, li2023informative also believe that the high-confidence nodes convey overlapping information, causing information redundancy in pseudo-labeled nodes. The representativeness of a node is computed with the mutual information between the node and its neighborhood aggregation. Pseudo-labeled nodes are selected based on representativeness and confidence scores exceeding predefined thresholds, with less confident nodes assigned lower weights in the gradient descent process. In this context, unlike in active learning where uncertainty is solely utilized for node selection, uncertainty within self-training also indicates the degree of noise present in pseudo-labels.
In conclusion, the application of uncertainty in node selection tasks is vital. A majority of current methods compute uncertainty with single deterministic models, especially direct estimation for efficient iterative training. For non-i.i.d. graph structures, the node uncertainty is computed along with its neighborhood influence and is often combined with other metrics to enhance performance. Nonetheless, we argue that in both graph active learning and self-training tasks, where the model is updated on expanded label sets, the overall predictive uncertainty of nodes is consistently decreasing. However, in each iteration, uncalibrated quantification methods may suffer from this distribution shift, as discussed in Section 3.4. Furthermore, to use the structural information from graphs more accurately, we suggest considering the overall uncertainty of node sets instead of the linear sum of node uncertainties with influence.
@@ 4.2 Uncertainty-based Abnormality Detection
OOD detection. To begin with, out-of-distribution (OOD) detection focuses on identifying instances that diverge from the training data distribution (in-distribution). OOD detection involves identifying samples from unknown classes and preventing models from making wrong decisions on these samples with overconfidence. In some literature (wu2020openwgl, ), OOD detection is integrated with the task of classification for in-distribution data and formulated as an open-world classification. Open-world classifiers not only recognize predefined classes seen during training but also identify instances that belong to new and previously unseen classes. For OOD detection, wu2020openwgl maximize an entropy-based uncertainty loss for unlabeled nodes during training, such that the instances of unseen classes can have a low and balanced predictive probability over the already seen classes. During inference, if the maximal class probability of an instance is lower than a predefined threshold, the instance is regarded as OOD. In addition, zhao2020uncertainty propose to quantify various types of uncertainty with graph-based Dirichlet distribution, and point out that vacuity uncertainty, derived from a lack of evidence or knowledge, is the most effective in OOD detection among other uncertainties. Furthermore, considering the non-i.i.d. nature of graph data, stadler2021graph perform Bayesian posterior updates for predictions on interdependent nodes through diffusing Dirichlet parameters. Epistemic uncertainty without network effects is shown effective in detecting random OOD nodes with perturbed features, while epistemic uncertainty with network effects performs better in identifying nodes from the left-out class not present in the training data.
@@
Outlier detection. Apart from OOD detection, outlier detection (also known as anomaly detection) is another prevailing task that effectively leverages uncertainty in graph learning. Outlier detection refers to the identification of data that exhibit significantly different behaviors or characteristics compared to the majority of the graph. Different from the OOD data, outliers can also exist in the training data. ray2021learning directly estimate the predictive uncertainty with deviation from the distribution for interpretable anomalous behavior detection on multivariate time series. Notwithstanding conventional uncertainty quantification methods, liu2022pygod interpret the graph anomaly scores using a Bayesian approach (perini2020quantifying, ), which converts the anomaly score into the confidence level of the Gnn in its outlier prediction for each example. The key idea of this conversion is to align the likelihood of an example being an outlier with its position in anomaly score distribution.
@@
Misclassification detection. Moreover, uncertainty can also be utilized to identify instances with incorrect model prediction, known as misclassification detection. zhao2020uncertainty attempt various types of uncertainty, and concludes that dissonance (i.e., uncertainty due to conflicting evidence) is the most effective in misclassification detection. Besides detecting OOD nodes, stadler2021graph also utilize their proposed method to detect misclassified nodes. In most experiments, aleatoric uncertainty achieved the best performance, rather than epistemic uncertainty in OOD detection. This observation supports the necessity to separate and utilize task-specific sources of uncertainty.
@@
In summary, recent studies emphasize the importance of uncertainty in graph-based anomaly detection, although there is inconsistency in the type of uncertainty utilized for the same purpose. For instance, both predictive uncertainty (wu2020openwgl, ), epistemic uncertainty (stadler2021graph, ), and vacuity uncertainty (zhao2020uncertainty, ) have been applied to detect OOD nodes. There is a lack of a systematic framework for selecting quantification methods and applying estimated uncertainty in abnormality detection. Establishing such a framework would be advantageous for both the theoretical exploration and the practical application of abnormality detection.

### v2

@@ 3.3 Other Methods
Different from the Bayesian Gnns that perturb parameters around a single optimal model, ensemble models derive the final prediction based on the predictions from a set of independent models with different parameters or even with different model structures. Ensemble Gnns have been shown to provide more accurate prediction and reliable estimation than the single models (Bazhenov et al., 2022; Wang et al., 2023; Gawlikowski et al., 2023). Meanwhile, ensemble Gnns offer an intuitive means of expressing predictive uncertainty by assessing the diversity among the predictions of the ensemble members. AU and EU could also be differentiated similarly as in Bayesian Gnn methods. Despite their wide applications, most Gnn ensemble models follow the deep ensemble framework (Lakshminarayanan et al., 2017) without specific modifications for both graphs and uncertainty quantification purposes. Compared with the two families above, ensemble methods are less used in Gnns mainly due to their high computational cost.
Test-time data augmentation, another type of uncertainty quantification method, has been widely used in other fields, e.g., computer vision (Ayhan & Berens, 2022). It involves augmenting multiple samples for each test sample and applying the model to all samples to compute the distribution of predictions. Then, AU is estimated with the entropy of the model prediction distribution for a test sample and its augmentations, and EU can also be estimated in the same way as Bayesian Gnn methods. However, in Gnns, test-time augmentation has only been utilized for other purposes. Jin et al. (2022) first propose to modify the node feature and the graph topology at the test time for better OOD generalization. Similarly, Ju et al. (2024) use test time augmentation to mitigate the bias of Gnn towards achieving better performance on high-degree nodes. However, test-time augmentation has not yet been explicitly applied for graph uncertainty quantification and thus, is not included in our primary framework. Nevertheless, its potential adaptation for graph uncertainty quantification warrants exploration. Given its post hoc nature, it requires no additional data and shows better computational efficiency during training compared to ensemble and Bayesian Gnns.
1.  Wang et al. (2021); Cai et al. (2017); Guo et al. (2017); Hsu et al. (2022b); Xu et al. (2023) 2.  Zhao et al. (2020); Stadler et al. (2021); Bazhenov et al. (2023) 3.  Kang et al. (2022); Efron (1982); Huang et al. (2023a); Zargarbashi et al. (2023) 4.  Zhang et al. (2019b); Hasanzadeh et al. (2020); Rong et al. (2019); Cha et al. (2023); Wu et al. (2021) 5.  Bazhenov et al. (2022); Wang et al. (2023); Gawlikowski et al. (2023)
@@
3.4 Evaluating Quantification Models
@@
We discuss the evaluation measures to assess the quality of Gnn uncertainty quantification models. As in Figure 2, evaluations are not limited to a specific type of quantification model but to each targeted uncertainty source. This is because a model can quantify multiple sources of uncertainty, and models can only be evaluated regarding a specific uncertainty source. Given the challenge of acquiring ground truth uncertainty, various metrics have been proposed, and there is no consensus on a unified metric (Lakshminarayanan et al., 2017; Gawlikowski et al., 2023).
Predictive (Total) uncertainty. As discussed above (Liu et al., 2022a; Zhao et al., 2020), maximum softmax probability (referred to as confidence in this paper) and entropy are widely applied to quantify the predictive uncertainty in the classification problems. They are also utilized as uncertainty measures of predictive uncertainty Smith & Gal (2018). Notably, entropy is often used as a measure of aleatoric uncertainty when applied on labeled dataset, e.g., in decision tree. However, entropy in Gnns is calculated on the predicted probability of each sample. We believe it should be considered to represent predictive uncertainty, as opposed to aleatoric uncertainty evaluation metrics as in Gawlikowski et al. (2023). As mentioned in Section 3.1, these two metrics are vulnerable to distribution shifts and need calibration.
In regression problems, the mean width of prediction intervals (MWPI) (Kang et al., 2022; Huang et al., 2023a) can be used for evaluation. The coverage rate—the portion of ground-truth labels falling into these intervals—is verified to ensure the validity of this metric. Notably, efficiency is the same notion as MWPI in conformal regression, while in conformal classification, it refers to the average prediction set size under given coverage rate. In some works (Zargarbashi et al., 2023), this metric is accompanied by singleton hit rate, i.e., the prediction accuracy of prediction sets of size one. This joint use of both metrics satisfies the need for both prediction quality and uncertainty quantification. However, MWPI estimates the predictive uncertainty over a set of data but not on individual samples.
@@
Model & Aleatoric uncertainty. The model uncertainty is commonly measured through mutual information (MI) to evaluate quantification models. MI is minimal when information in prediction does not increase with additional model knowledge. It is calculated as the difference between the entropy of the expected distribution (predictive uncertainty) and the expected entropy (aleatoric uncertainty) (Malinin & Gales, 2018; Zhao et al., 2020). Notably, MI is also referred to as information gain in some papers (Liu et al., 2022a). For Bayesian and ensemble Gnns, the predictive variance of sampled outputs is another option to interpret model uncertainty (Zhang et al., 2019a; Gal & Ghahramani, 2016). Furthermore, employed as one of the optimization objectives in some quantification methods (Liu et al., 2020; Zhao et al., 2020), the expected Kullback-Leibler (KL) divergence between approximated distribution and the posterior distribution can also evaluate model uncertainty. Given no distributional uncertainty, model uncertainty equals epistemic uncertainty. In this case, aleatoric uncertainty, namely data uncertainty, can be measured as the difference between the total uncertainty and model uncertainty; otherwise, the metric will be no longer meaningful.
@@
Distributional uncertainty. It is believed that quantification of distributional uncertainty will be able to separate in-distribution (ID) and out-of-distribution (OOD) samples Malinin & Gales (2018). A popular approach to measure total distributional uncertainty is through classifying test samples. Samples with distributional uncertainty surpassing a given threshold are categorized as OOD samples, while those below the threshold are considered ID samples. In this, conventional classification metrics, e.g., Receiver Operating Characteristic (ROC) curve and Area Under Receiver Operating Curve (AUROC) (Zhao et al., 2020; Song & Wang, 2022; Gui et al., 2022), can be applied to measure the distributional uncertainty. However, this evaluation depends on the selection of test samples and threshold.
In conclusion, the effectiveness of existing metrics remains under exploration. As an alternative, many studies directly employ specific downstream tasks for evaluation to obtain reliable and straightforward measures. Nonetheless, depending on demands under different circumstances, we recommend diverse methods quantification methods according to their properties, as shown in Figure 3. First, it is crucial to identify the specific source of uncertainty being targeted. To our knowledge, only Bayesian-based estimation methods explicitly quantify distributional uncertainty. Subsequently, one must consider the necessity of separating aleatoric and epistemic uncertainty. If this separation is deemed necessary, two categories of methods are available: ensemble models, which consist of different member models, and single models with parameters treated as random variables. For computational efficiency, we recommend single models with parameters as random variables. However, for improved prediction performance and robustness across different distributions, ensemble models are preferable. In some cases, the separation of aleatoric and epistemic uncertainty does not yield significant benefits. For practitioners seeking highly computationally efficient methods for individual examples (e.g., nodes or links), direct estimation is adequate. Conversely, for a distribution-free and statistically guaranteed estimation of the overall test population, Frequentist-based estimation is recommended.
Figure 3: Flowchart illustrating recommended quantification methods for various conditions. Diamond shapes represent conditions, while rectangular shapes indicate the recommended methods.
@@ 4.1 Uncertainty-based Node Selection
Active learning. In graph active learning tasks, uncertainty and representativeness (diversity) are prevalent criteria for node selection  (Ren et al., 2021). Uncertainty is often measured by the entropy of the prediction probabilities on the nodes. Cai et al. (2017) first propose a graph active learning framework, which uses a combination of entropy-measured uncertainty, information density, and node centrality to select nodes. Zhang et al. (2021a) estimate an ensembled uncertainty using the variation of prediction generated by committee models, which is more robust and reliable than entropy as a single deterministic model. The method selects top $k$ nodes to maximize the effective reception field using the topological information of the graph, excluding nodes that are less uncertain and less informative. Unlike the previous two methods, which select the most uncertain nodes independently, Zhang et al. (2021b) consider label propagation and select a subset of nodes with maximized uncertainty (entropy) reduction to include diversity. Labeled nodes can propagate their label information based on the graph distance and the node feature similarity, and then reduce the uncertainty of adjacent unlabeled nodes. Additionally, in a multi-task manner, Chang et al. (2024) estimate the confidence difference between the node classification task and anomaly detection task for node selection in graph active learning. Notably, in most cases, graph active learning methods that select nodes with only uncertainty can hardly outperform the ones with a combination of uncertainty and other measures.
@@
To achieve dynamic and adaptive node labeling, reinforcement learning (RL) has been applied to graph active learning. Gao et al. (2018) uses the combination of the measures in Cai et al. (2017) to select nodes, but dynamically adjusts the combination weights based on a multi-armed bandit framework. Instead of measuring nodes independently, Hu et al. (2020) formulates the active learning problem as a sequential decision process on the graph considering the interconnections between nodes. Uncertainty is still measured with entropy but concatenated with other features, e.g., the divergence between a node’s predicted label distribution and its neighbor’s, to form the state representation of each node. The reward is the model performance on hold-out data updated with a trajectory of selected nodes. However, the RL-based active learning methods are significantly time-consuming to train.
@@
Self-training. In graph self-training, i.e., pseudo-labeling, Li et al. (2018) propose one of the initial works that expand the labeled set with top $K$ confident nodes, quantified through maximal softmax probabilities. Considering that Gnn is ineffective in propagating label information under the few-label setting, Sun et al. (2020) iteratively select and label the top $K$ confident nodes. Drawing upon the concept of ensembling to robustly select and label nodes, Yang et al. (2021) employs multiple diverse Gnns. Nodes are selected with confidence surpassing a pre-determined threshold and ensuring consistent prediction for labels across all models. Wang et al. (2021) is the first work to point out that the under-confidence of Gnns leads to biased uncertainty estimation, deteriorating the performance of uncertainty-based self-training node selection. With the calibrated confidence, they improve the performance of the self-training method introduced by Li et al. (2018). In consideration of graph structure information, Zhao et al. (2021) propose to compute entropy for each node, and add an entropy-aggregation layer to Gnn for uncertainty estimation.
@@
Similar to graph active learning, recent self-training methods (Liu et al., 2022a; Li et al., 2023b) consider combining uncertainty with representativeness, but from different perspectives. Liu et al. (2022a) believe the common self-training practice that only selects high-confidence nodes leads to the distribution shift from the original labeled node distribution to the expanded node distribution. Therefore, they follow the common practice that selects high-confidence nodes as pseudo-labels but decreases the weight of the nodes with low information gain in the training loss. Interestingly, the dropout technique employed to compute information gain is often interpreted as the model uncertainty, as introduced in Section 3.4. Similarly, Li et al. (2023b) also rely on the fact that the high-confidence nodes convey overlapping information, causing information redundancy in pseudo-labeled nodes. The representativeness of a node is computed with the mutual information between the node and its neighborhood aggregation. Pseudo-labeled nodes are selected based on the representativeness and confidence scores exceeding predefined thresholds, with less confident nodes assigned lower weights in the gradient descent process. Further, Wang et al. (2024) reformulates the goal of selecting pseudo-labeled nodes as to reduce the uncertainty over all unlabeled nodes, and consider node selection in a combinatorial manner. This method can optimize the student model with a more global approach and reduces the correlation of selected nodes, improving representativeness and effectiveness of node selection in graph self-training.
Quantifying uncertainty accurately during the iteration process remains a significant challenge in both graph active learning and self-training tasks. One key issue is that the distribution of newly added pseudo-labeled data often diverges from the original training data distribution, making uncertainty estimates derived from the existing training data uncalibrated and noisy. In self-training, where teacher models frequently make inaccurate predictions, pseudo-labels can further introduce noise into uncertainty estimation. While limiting predictions to labeled data can reduce noise in some scenarios, it fails to capture the evolving uncertainty of updated models throughout the self-training iterations. Graph calibration models, as discussed in Section 3.1, offer a potential solution to this challenge by recalibrating uncertainty estimates. However, these models require a sufficient amount of labeled calibration data and typically involve longer training times. Moreover, in graph active learning, uncertainty serves primarily for node selection. By contrast, in self-training, it also represents the degree of noise in pseudo-labels. Unfortunately, inaccurate pseudo-labels can undermine the effective use of uncertainty, as even predictions with low uncertainty may contain incorrect labels, potentially leading to model misdirection during retraining.
In summary, uncertainty plays a crucial role in node selection tasks. Most of the existing methods rely on single deterministic models for estimating uncertainty directly, favoring computational efficiency in iterative training. For non-i.i.d. graph structures, node uncertainty is calculated alongside the influence of neighboring nodes and is often combined with other performance-enhancing metrics. Despite its widespread application, researchers should be cautious when estimating and applying uncertainty, particularly in graph self-training scenarios.
@@ 4.2 Uncertainty-based Abnormality Detection
OOD detection. To begin with, out-of-distribution (OOD) detection focuses on identifying instances that diverge from the training data distribution (in-distribution). OOD detection involves identifying samples from unknown classes and preventing models from making wrong decisions with overconfidence. In some literature (Wu et al., 2020a), OOD detection is integrated with the task of classification for in-distribution data and formulated as an open-world classification. Open-world classifiers recognize predefined classes seen during training and identify instances from new classes. For OOD detection, Wu et al. (2020a) maximize an entropy-based uncertainty loss for unlabeled nodes during training, such that the instances of unseen classes can have a low and balanced predictive probability over the already seen classes. During inference, if the maximal class probability of an instance is lower than a predefined threshold, the instance is regarded as OOD. In addition, Zhao et al. (2020) propose to quantify various types of uncertainty with graph-based Dirichlet distribution, and point out that vacuity uncertainty, derived from a lack of evidence or knowledge, is the most effective in OOD detection among other uncertainties. Furthermore, considering the non-i.i.d. nature of graph data, Stadler et al. (2021) perform Bayesian posterior updates for predictions on interdependent nodes through diffusing Dirichlet parameters. Epistemic uncertainty without network effects is shown effective in detecting random OOD nodes with perturbed features, while epistemic uncertainty with network effects performs better in identifying nodes from the left-out class not present in the training data.
@@
Outlier detection. Apart from OOD detection, outlier detection (also known as anomaly detection) is another prevailing task that effectively leverages uncertainty in graph learning. Outlier detection refers to the identification of data that exhibit significantly different behaviors or characteristics compared to the majority of the graph. Different from the OOD data, outliers can also exist in the training data. Ray et al. (2021) directly estimate the predictive uncertainty with deviation from the distribution for interpretable anomalous behavior detection on multivariate time series. Notwithstanding conventional uncertainty quantification methods, Liu et al. (2022c) interpret the graph anomaly scores using a Bayesian approach (Perini et al., 2020), which converts the anomaly score into the confidence level of the Gnn in its outlier prediction for each example. The key idea of this conversion is to align the likelihood of an example being an outlier with its position in anomaly score distribution.
@@
Misclassification detection. Moreover, uncertainty can also be utilized to identify instances with incorrect model prediction, known as misclassification detection. Zhao et al. (2020) attempt various types of uncertainty, and concludes that dissonance (i.e., uncertainty due to conflicting evidence) is the most effective in misclassification detection. Besides detecting OOD nodes, Stadler et al. (2021) also utilize their proposed method to detect misclassified nodes. In most experiments, aleatoric uncertainty achieved the best performance, rather than epistemic uncertainty in OOD detection. This observation supports the necessity to separate and utilize task-specific sources of uncertainty.
@@
In summary, recent studies emphasize the importance of uncertainty in graph-based anomaly detection, although there is inconsistency in the type of uncertainty utilized for the same purpose. For instance, both predictive uncertainty (Wu et al., 2020a), epistemic uncertainty (Stadler et al., 2021), and vacuity uncertainty (Zhao et al., 2020) have been applied to detect OOD nodes. There is a lack of a systematic framework for selecting quantification methods and applying estimated uncertainty in abnormality detection. Establishing such a framework would be advantageous for both the theoretical exploration and the practical application of abnormality detection.